OpenStudy (itsbribro):

Acceleration is defined as the rate of change for which of the following? A. time B. velocity <---- C. position D. displacement

OpenStudy (sreekar369):

Acceleration is defined as the rate of change velocity

OpenStudy (sreekar369):

@itsbribro

OpenStudy (daniel_chernioglo):

I think your correct but let me double check

OpenStudy (daniel_chernioglo):

nvm i cant find it, but yes you are correct

OpenStudy (itsbribro):

thank you

OpenStudy (itsbribro):

ok

OpenStudy (itsbribro):

haha okay thank you can you help me with some more?

OpenStudy (sreekar369):

thq u for the medal @its

OpenStudy (sreekar369):

@itsbribro

OpenStudy (itsbribro):

np

OpenStudy (daniel_chernioglo):

ya i can try?

OpenStudy (sreekar369):

i will help

OpenStudy (itsbribro):

What is the acceleration of an object if it goes from a velocity of +25 m/s to rest in 5.0 s? A. –5 m/s2 B. +5 m/s2 C. –25 m/s2 D. +25 m/s2

OpenStudy (sreekar369):

is option a is -5

OpenStudy (itsbribro):

the question marks are negatives

OpenStudy (sreekar369):

if it is answer is -5

OpenStudy (itsbribro):

okay

OpenStudy (itsbribro):

What is the average acceleration of a jet ski that goes from rest to 10 m/s west in 4 s? Take the direction west to be positive, and east to be negative. A. 2.5 m/s2 <---- B. -2.5 m/s2 C. 0.4 m/s2 D. -0.4 m/s2

OpenStudy (sreekar369):

a=v-u/t as the object came to rest,v=0 a=-u/t a=-25/5 so,-5 is answer

OpenStudy (sreekar369):

this is for before question

OpenStudy (sreekar369):

option a

OpenStudy (itsbribro):

ok the explination thank you

OpenStudy (itsbribro):

*explanation

OpenStudy (sreekar369):

same formula for this question but now it came from rest so take "u" as 0 and substitute the given values

OpenStudy (itsbribro):

thx @sreekar369

OpenStudy (sreekar369):

if u didnt got it say to me

OpenStudy (sreekar369):

no prblm @itsbribro

OpenStudy (itsbribro):

thank you so much i got 100% i had 5 more questions after that :)

OpenStudy (sreekar369):

ok close this and ask a new question